In a large hollow cylinder, roller-coaster rider

can drive the car along the wall of the cylinder.

What minimum speed must be the car

without sliding downward?


1
Expert's answer
2021-11-07T19:25:09-0500

The minimum speed must provide enough potential energy to reach the top and remain upside down:


"mg=ma_c,\\\space\\
mg=m\frac{v^2}R,\\\space\\
v=\sqrt{gR}."